### v1

# Unveiling Linguistic Regions in Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated considerable cross-lingual alignment and generalization ability. Current research primarily focuses on improving LLMs’ cross-lingual generalization capabilities. However, there is still a lack of research on the intrinsic mechanisms of how LLMs achieve cross-lingual alignment. From the perspective of region partitioning, this paper conducts several investigations on the linguistic competence of LLMs. We discover a core region in LLMs that corresponds to linguistic competence, accounting for approximately 1% of the total model parameters. Removing this core region by setting parameters to zero results in a significant performance decrease across 30 different languages. Furthermore, this core region exhibits significant dimensional dependency, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence. Moreover, we discover that distinct regions exist for different monolingual families, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages. Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common occurrence observed during further pre-training of LLMs. Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence.
Unveiling Linguistic Regions in Large Language Models
[…]
1 Introduction
Figure 1: Three main findings of our experiments: (1) Identification of core language regions within the LLMs, where removals lead to linguistic competence loss; (2) Discovery of monolingual regions, where removals cause significant proficiency loss in specific languages; (3) Optimization of freezing core regions during further pre-training decelerates language forgetting.
Over the years, the field of Natural Language Processing (NLP) has been at the forefront of understanding the core principles of intelligence. The emergence of Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023), and LLaMA 2 (Touvron et al., 2023), showcases a significant breakthrough. Thanks to unparalleled scales of model architecture and the vastness of training data, these LLMs now exhibit exceptional linguistic competence and can execute complex tasks requiring abstract knowledge (Dong et al., 2023) and reasoning (Cobbe et al., 2021).
Previous research has revealed that LLMs naturally capture cross-linguistic similarities in their representation space, facilitating zero-shot cross-lingual transfer (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023). The model is fine-tuned on one language, enabling the acquisition of comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), and exhibits the phenomenon of code-switching when generating context, switching between languages within a single utterance (Khanuja et al., 2020; Zhao et al., 2024). Attempts to improve LLMs’ cross-lingual generalization abilities have been successful through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), aligning languages compulsorily (Sherborne and Lapata, 2022; Shaham et al., 2024) and utilizing in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023). However, a detailed research of underlying the internal mechanisms of how LLMs possess cross-linguistic alignment capability remains elusive.
To delve deeper into the intrinsic mechanisms of LLMs’ linguistic competence, this paper focus on the LLMs’ parameter importance and investigate the linguistic regions of LLMs based on 30 distinct languages’ performance, with the purpose of figuring out the following questions:
Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual generalization? Through pre-training across six languages and measurement of parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters. As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently lead to a significant decline in performance across 30 test languages (Section 3.2).
Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependency. In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3).
Q2: Beyond the core multilingual region within LLMs , do distinct monolingual (or monolingual family) regions exist that specifically influence individual languages? While LLMs possess strong multilingual capabilities, we discover that each individual language (or language family) encompasses independent regions within the LLMs. As shown in the middle of Figure 1, the analysis of the Russian sentences identifies a particular linguistic region that likewise exerts influence both on the Russian and Ukrainian language, both of which belong to the Slavic group (Section 3.4).
Q3: If and how core linguistic regions affect further pre-training, how to utilize it to optimize further pre-training? After pre-training, core linguistic parameter regions of the LLMs are established for multilingual alignment. Notable shifts in these regions potentially lead to a decline in model lingual capabilities. Our findings reveal that freezing this core region can mitigate the issue of catastrophic forgetting (McCloskey and Cohen, 1989; Kemker et al., 2018), a common occurrence observed during further pre-training of LLMs. As shown at the bottom of Figure 1, we investigate the impact of selectively freezing 5% key parameters of all parameters fine-tuning during further pre-training, compared to the full-scale fine-tuning technique. Findings indicate that this method facilitated comparable or even more efficient learning of the target language, while concurrently decelerating the rate of language attrition for previously learned languages (Section 5). Significantly, our methodology is compatible with the data-replay techniques (Robins, 1995), with no necessity for integrating extra components into the model. Unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), our approach restricts to a minimal core region in LLMs.
The main contributions of our work are summarized as follows:
  • We discover that LLMs possess a core linguistic region, and removing this region (setting parameters to zero) results in a significant loss of the model’s linguistic capabilities. Furthermore, perturbations to specific dimensions or even a single parameter can lead to a substantial decline in the model’s linguistic abilities.
  • We observe that distinct monolingual regions exist in LLMs for different languages (or language families). Removing a specific monolingual region causes a significant deterioration in the linguistic capabilities within that language (or language family).
  • We perform further pre-training for specific languages within the core linguistic region of LLMs frozen, achieving comparable or even superior performance in the target language while mitigating catastrophic forgetting in non-target languages.
2 Background and Metric
2.1 Model Pre-training
Pre-training is a crucial process by which LLMs acquire linguistic competence and gain general knowledge about the real world. Formally, given a large corpus $\mathcal{D}$, the training objective is to find the optimal $\theta$, minimize the following loss $\mathcal{L}$:
$\mathcal{L}{(\mathcal{D},\theta)}=\sum_{x\in\mathcal{D}}\sum_{i}\log p_{\theta}(x_{i}|x_{1},...,x_{i-1}),$ (1)
where $x=\{x_{1},...,x_{n}\}$ denotes an input token sequence and $\theta$ denotes parameters of the model.
2.2 Parameter Importance
Drawing upon the observations of linguistic alignments, we propose that particular parameters regions within the model exert significant influence on its inherent language alignment capabilities. Evaluating parameter sensitivity is a crucial metric for determining the significance of parameters in model pruning (Sanh et al., 2020; Liang et al., 2021; Zhang et al., 2022). If removing a parameter (zero-out) significantly affects the loss, the model is sensitive to it. More specifically, given a large corpus $\mathcal{D}$ and $\theta=[\theta_{1},\theta_{2},\ldots,\theta_{d}]\in\mathbb{R}^{d}$ as the parameters of a model, with each $\theta_{j}\in\mathbb{R}$ denoting the $j$-th parameter, the training objective is to minimize loss $\mathcal{L}{(\mathcal{D},\theta)}$ (defined in 2.1). The importance of each $\theta$ is denoted as $\mathcal{I}(\theta)\in\mathbb{R}^{d}$, where its $j$-th index $\mathcal{I}_{j}(\theta)$ signifies the importance for $\theta_{j}$.
Under an independent and identically distributed data (i.i.d.) assumption, the importance of a parameter $\mathcal{I}_{j}(\theta)$ is measured by the increase in prediction loss when it is removed, calculated as the absolute difference between prediction losses with and without the parameter($\theta_{j}$):
$\mathcal{I}_{j}(\theta)=\left|\mathcal{L}{(\mathcal{D},\theta)}-\mathcal{L}{(\mathcal{D},\theta|\theta_{j}=0)}\right|.$ (2)
Calculating $\mathcal{I}_{j}(\theta)$ for each parameter, as outlined in 2, is computationally expensive because it involves $d$ distinct versions of the network computing, for each removed parameter. This becomes particularly challenging as the number of model parameters, $d$, grows to hundreds of billions. However, according to the Taylor expansion formula for $\mathcal{L}$ at $\theta_{j}=0$:
$\quad\mathcal{L}(\mathcal{D},\theta)=\mathcal{L}(\mathcal{D},\theta|\theta_{j}=0)$ (3)
$+\frac{\partial\mathcal{L}}{\partial\theta_{j}}(\theta_{j}-0)+\frac{1}{2!}\frac{\partial^{2}\mathcal{L}}{\partial\theta_{j}^{2}}(\theta_{j}-0)^{2}+\cdots,$
we can estimate $I_{j}(\theta)$ with its first-order Taylor expansion, eliminating the requirement for $d$ distinct networks computation:
$\mathcal{I}_{j}(\theta)\approx\left|g_{j}\theta_{j}\right|,$ (4)
where $g_{j}=\frac{\partial\mathcal{L}}{\partial\theta_{j}}$ are elements of the parameter gradient $g$, and the importance is easily calculated since the gradient $g$ can be obtained from backpropagation.
3 Experiments
3.1 Experimental Setup
To localize the functional regions corresponding to linguistic competence within LLMs and analyze their nature, we perform language further pre-training (next token prediction) on various languages and observe the relationship between internal parameter removal and external output quality. We utilize LLaMA-2-7B/13B (Touvron et al., 2023) as our model instance, as it stands out as one of the most notable state-of-the-art open-source LLMs in current academia.
Our experimental dataset comprises materials from Chinese platforms like Zhihu and Wechat, English sources from Arxiv and Falcon, and a corpus including books from 28 languages, totaling 30 languages in all. Six languages, namely Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese, are chosen for language further pre-training and region localization, with $100,000$ samples for each (distinct from the samples in the test set). All 30 languages are employed for model testing and functional region analysis, with the specific languages and token count detailed in A. We use perplexity (PPL) as the criterion for evaluating the linguistic competence of a language model.
Languages LLaMA-2 3% Removal
Base Top Bottom Random
Arabic 6.771 127208.250 6.772 7.895
6.261 102254.758 6.316 7.112
Chinese 8.652 295355.5 8.565 9.837
7.838 84086.906 7.806 8.619
Italian 14.859 58908.879 14.860 17.341
13.694 47375.844 13.730 15.207
Japanese 10.888 322031.406 10.896 12.535
10.072 75236.031 10.137 11.661
Korean 4.965 125345.359 4.967 5.649
4.724 90768.844 4.743 5.241
Persian 6.509 81959.719 6.511 7.628
6.205 92201.812 6.229 7.009
Portuguese 15.318 47763.059 15.319 17.297
13.667 51498.402 13.982 15.376
Russian 12.062 170776.750 12.064 13.728
11.048 112574.609 10.948 11.757
Spanish 17.079 51940.859 17.082 18.98
16.351 54005.891 16.138 17.292
Ukrainian 9.409 120719.938 9.409 10.875
8.295 116287.305 8.297 9.076
Vietnamese 5.824 40126.527 5.824 6.614
5.471 42336.426 5.437 5.995
Table 1: LLaMA-2 perplexity on 11 languages with 3% removal ratio. The 13B model is gray-filled while the 7B model is unfilled. ‘Top’ and ‘Bottom’ respectively indicate the $N$ parameters with the highest and lowest cumulative $\mathcal{I}^{*}_{j}(\theta)$ during the further pre-training across the six languages. ‘Random’ denotes the randomly selecting $N$ while ‘Base’ represents no removal. Here, $N$ equals 3% of the total number in each parameter matrix.
3.2 Core Linguistic Competence Region
In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify core parameter regions associated with linguistic competence. Specifically, according to Equation 4, we cumulatively compute *$\mathcal{I}^{*}(\theta)=\Sigma\mathcal{I}(\theta)$ values across six different languages’ training, positing that the set of parameters exhibiting maximal importance score *$\mathcal{I}^{*}(\theta)$ during the language further pre-training may have a strong correlation with the model’s linguistic competence, and we provide both logical and empirical evidence to support this hypothesis.
Logical Evidence
The phenomenon of code-switching suggests that the LLMs can align languages and may possess core linguistic regions. As discussed in Section 2.2, if a parameter $\theta_{j}$ is crucial for the LLMs’ core linguistic competence, the model should be sensitive to $\theta_{j}$, shown by a significant increase on the loss $\mathcal{L}$ when $\theta_{j}$ is removed, severely impairing the LLMs’ linguistic performance. Conversely, other parameters impact rarely on core linguistic capabilities.
Empirical Evidence 1
Table 1 illustrates that even a 3% removal on the ‘Top’ region leads to a substantial increase in perplexity (PPL), reaching over $40,000$ across 11 languages, indicating a complete loss of linguistic competence. In contrast, removing the ‘Bottom’ region is comparable to non-removal ‘Base’ in model PPL, and a ‘Random’ removal of equal magnitude has no significant impact on the model’s linguistic competence. Moreover, refer to Appendix B, additional experiments with reducing training samples to $10,000$ or adjusted the region selection ratio to 1% and 5% yield consistent findings: removing the ‘Top’ region deprives LLaMA-2 of its capability across all 30 languages. This suggests the model’s linguistic competence is directly influenced by the ‘Top’ region, while removing the ‘Bottom’ and ‘Random’ region don’t have a significant direct impact on language capabilities. See Appendix B for evaluations on 30 languages and further experiments.
Figure 2: Visualization of the linguistic competence region (the ‘Top’ 5% region). The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the Top region.
Testing Dataset (Language) # Training Samples (Chinese) Removal Ratio $=1\%$
Top & Freeze Bottom & Freeze Top & Unfreeze
Wechat (Chinese) 0K 254772480 6.452 254772480
2K 674.076 6.052 6.05
5K 292.499 6.053 6.058
10K 116.859 6.305 6.303
20K 20.722 6.556 6.559
50K 9.129 6.18 6.175
200K 6.246 5.581 5.604
Falcon (English) 0K 4244070 14.02 4244070
2K 158431.282 14.507 14.445
5K 343498 15.732 15.415
10K 175567.219 15.878 15.875
20K 32505.828 18.689 18.952
50K 12455.038 29.029 31.583
200K 5301.527 488.429 448.804
Table 2: Removing-freezing analysis at $1\%$ removal ratio in different regions of LLaMA-2-7B. ‘Top/Bottom’ denotes the removal region, while ‘Freeze/Unfreeze’ indicates whether the corresponding region is frozen after removal.
Empirical Evidence 2
In the experiment corresponding to Table 2, we initially zero out various regions within LLaMA. Consistent with the findings from Table 1, removing the ‘Top’ region leads to a loss of linguistic competence, whereas the ‘Bottom’ region don’t. However, in this experiment, we sought to ascertain if LLaMA could reacquire its lost cross-lingual generalization competence. Thus, we train on different amounts of Chinese Zhihu corpus and evaluate on Chinese Wechat and English Falcon corpora. The results indicate that unlike the ‘Bottom’ region, if the ‘Top’ region is removed and frozen, the model have to relearn basic language rules in other regions based on the provided Chinese Zhihu corpus, but these rules are inherently biased towards Chinese. Consequently, while its proficiency in Chinese is restored, the English perplexity remains high ($5301.527$). If the ‘Top’ region is removed but not frozen, the model can rebuild its linguistic competence in-place. As its proficiency in Chinese is restored, so is its proficiency in English. This implies that the ‘Top’ region encodes generalizable fundamental linguistic competence. When ‘Top’ region is zeroed-out and frozen, other regions significantly adapt to regain core linguistic competence. Similar conclusions can be obtained with an expanded removal ratio 5%. For further details, see Appendix C.
Model Size # Training Samples $N_{d}$ Attn.o(row), Attn.k/q/v+FFN.down(column)
Top Middle Bottom Random
7B 100K $1$ 848.326 6.447 6.447 6.48
100K $3$ 72594.445 6.455 6.458 6.487
100K $5$ 48001.992 6.461 6.463 6.495
100K $10$ 62759.516 6.478 6.48 6.529
13B 100K $1$ 5218.1 5.857 5.857 5.856
100K $3$ 37344.078 5.863 5.858 5.985
100K $5$ 41840.613 5.867 5.86 5.89
100K $10$ 465740.125 5.879 5.869 6.992
13B 10K $1$ 23120.977 5.859 5.856 5.865
10K $3$ 28816.867 5.862 5.86 5.875
10K $5$ 73268.289 5.866 5.862 5.878
10K $10$ 592922.25 5.879 5.871 5.993
Table 3: Perplexity of LLaMA-2 after removing certain dimensions in the Attention and Feedforward layers. Here, $N_{d}$ denotes the number of dimensions to remove, ’Top’, ’Middle’, and ’Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training. ’Random’ denotes an equivalent number of dimensions chosen randomly for comparison.
3.3 Dimensional Dependence
To provide a more intuitive revelation of the spatial distribution characteristics of the linguistic competence region within the model, we visualize the ‘Top’ region. As shown in Figure 2, whether in the attention mechanism layer or the feed-forward layer, the linguistic region displays a distinct concentration in both the rows and columns of the matrices. Such distribution features seem to imply that the model’s linguistic competence is concentrated in specific dimensions.
Structured Removal
Instead of discretely removing different unstructured parameters, we selectively remove structured certain rows or columns for each matrix, especially those dimensions encompassing a significant number of ‘Top’ region parameters, termed as ‘Top’ dimensions. As illustrated in Table 3, we attempt to remove the columns of FFN.down and Attn.k/q/v, as well as the rows of Attn.o. The results indicate that removing just these ‘Top’ dimensions leads to a substantial decline in the model’s linguistic competence. However, disturbances to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects. Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in D.
Single Dimension Perturbation
Here, we explore whether a specific dimension significantly impact the model’s linguistic competence. As illustrated in Figure 3, we iterate through the key dimensions mentioned in Section 3.3, attempting to perturb the same dimension (random initialization) across all Transformer layers. The results indicate that the impact of dimensions 2100 and 4743 on the LLaMA-2-13B substantially surpassed other dimensions, even when compared to the other three in the ‘Top5’ dimensions. In contrast, perturbing two randomly selected dimensions, 2800 and 4200, yield linguistic performance almost indistinguishable from the unperturbed state.
Figure 3: Perplexity of the LLaMA-2-13B when perturbing the same single dimension (Att.O and FFN.down matrices) across all layers. ‘Topk’ represents the top k dimensions that disrupt the model the most. ‘Random selected’ refers to a randomly chosen dimension. ‘Original’ indicates that no dimensions are disrupted.
Single Parameter Perturbation
We discover that even a slight modification to a single parameter in a model with over 13 billion parameters can lead to a significant decline in its output quality. In Table 4, merely resetting the $2100$-th parameter in the ‘Input_LayerNorm’ module of the $1$-st layer to its initial value causes LLaMA-2-13B’s PPL value to skyrocket from 5.865 to 83224.078. If this weight parameter is multiplied by $10$, the PPL value also rises to $4363.462$. This suggests that even minor changes to a single parameter can cause the model to lose nearly all of its linguistic competence. However, randomly altering the parameters at dimensions 2800 and 4200 doesn’t noticeably impact the model. For more details, refer to Appendix E.
Output Under Perturbations
To visually illustrate the impact of the linguistic competence region on the model’s output quality, we use "Fudan University is located in" as a premise and observe the model’s outputs under different parameter perturbations. The results are shown in Figure 4. Compared to randomly selected dimension 4200-th, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
3.4 Monolingual Family Region
In this section, we wonder if LLMs possess distinct regions within an individual language (or language family). Unlike the core linguistic regions described in Section 3.2, a monolingual family region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.
Perturbation Parameter Perplexity
- - 5.865
Reset 1 L1-N2100 83224.078
Reset 1 L1-N2800 5.860
Reset 1 L1-N4200 5.858
Mul 10 L1-N2100 4363.462
Mul 10 L1-N2800 5.859
Mul 10 L1-N4200 5.864
Table 4: Perplexity of LLaMA-2-13B on Chinese when perturbing a single weight parameter. Here, ‘Reset $1$’ represents resetting the parameter to $1$ (the initial value before pre-training), ‘Mul 10’ represents multiplying the parameter by $10$. ‘L1’ represents $1$-st layers. ‘N’ represents the ‘Input_LayerNorm’ module, followed by the perturbed dimension.
Region Localization
Different from Section 3.2, we initially identify and select the 1% ‘Top’ and ‘Bottom’ regions for each of the six languages (Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese) according to Equation 4, then deduplicate these regions. For the target language region, we exclude any regions that overlap with the ‘Top’ and ‘Bottom’ regions of the other five languages, aiming to eliminate the core regions and critical dimension corresponding to the model’s fundamental linguistic abilities. We denote $L$, $\mathcal{S}$ and $\mathcal{S}^{*}$ as the total set of six languages and the ‘Top/Bottom’ regions before and after deduplication, respectively. Language $l$’s own region $\mathcal{S}_{l}^{*}$ is computed as follows:
$\mathcal{S}_{l}^{*}=\mathcal{S}_{l}\,\,-\bigcup_{l^{{}^{\prime}}\in L\setminus\{l\}}\mathcal{S}_{l^{{}^{\prime}}}.$ (5)
Region Removal
Unlike removing core regions or dimensions in Section 3.2 and 3.3, we discover that removing monolingual regions will only significantly affect the ability of the target languages and their closely related languages with similar letter elements or sentence structure. For example, if we remove only the region $\mathcal{S}_{Russian}^{*}$ for Russian alone, selected from $10,000$ (10K) or $100,000$ (100K) samples respectively, as shown in Table 5, only Russian itself and Ukrainian have significant increases in PPL when removing ‘Top’ region. We speculate this to the fact that Russian and Ukrainian are relatively similar in terms of sentence structure and constituents, both belonging to the Slavic group. A similar phenomenon is observed if removals are changed to the regions for each of other five languages, see Appendix F for more details.
Figure 4: Comparison of linguistic competence. Expanding a single parameter to four times leads to error language competence in LLaMA-2-13B, a 13 billion-parameter LLM.
3.5 Further Pre-training Optimization
In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while preserving performance comparable to full-scale fine-tuning in target language proficiency. Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of $256$, a maximum token count of $512$, and learning rates ($lr$) and $5e{-5}$ or $5e{-6}$, employing perplexity (PPL) as the evaluation criterion.
Full-scale Model Fine-tuning
Traditional full-scale fine-tuning, when increasing the learning rate, enhances learning in the target language but aggravates forgetting in non-target languages. To counteract this forgetting, it is often essential to incorporate a portion of data from these other languages. As depicted on the left side of Figure 5 in line Blue, increasing $lr$ from $5e{-6}$(dotted blue line) to $5e{-5}$ (solid blue line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while accelerates the forgetting rate of the non-target languages (English and Chinese) simultaneously, shown in the middle and right side.
Freeze Core Regions Fine-tuning
We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment. To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 5. At a $lr$ of $5e{-6}$ (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal. However, when the $lr$ increases to $5e{-5}$ (solid line), freezing fine-tuning not only facilitates faster learning in the target language (achieving better performance in Arabic PPL: 3.557 vs. 3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs. 20.557 and 90.84 vs. 563.423 , respectively).
The potential reason for this phenomenon may lie in the preservation of the core regions within the cross-lingual alignment competence. Restricting the magnitude of updates in the core region’s parameters is a future strategy we intend to employ. Notably, unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), such approaches restricts to a minimal core region in LLMs, and can be implemented alongside blending previous data, retraining the entire network, or possibly only the final layers, without adding additional components to the model.
Languages Russian (10K) Russian (100K)
Base Top Bottom Top Bottom
Arabic 6.771 7.105 6.785 7.071 6.787
Chinese 8.562 8.927 8.593 8.878 8.599
Italian 14.859 16.155 14.931 16.274 14.935
Japanese 10.888 11.212 10.931 11.119 10.951
Korean 4.965 5.19 4.972 5.149 4.974
Persian 6.509 6.93 6.506 6.894 6.515
Portuguese 15.318 16.51 15.247 16.421 15.247
Russian 12.062 28.93 12.141 41.381 12.137
Spanish 17.079 18.07 17.224 17.894 17.211
Ukrainian 9.409 18.147 9.43 22.622 9.435
Vietnamese 5.824 6.086 5.872 6.079 5.873
Table 5: LLaMA-2-7B perplexity on 11 languages with a Russian region removal. Here, ‘Arabic’ and ‘Persian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged. Values with greater changes compared to the other regions’ removals are in bold.
Figure 5: Perplexity of LLaMA-2 across Arabic, English, and Chinese when training on 100,000 Arabic sentences. Blue represents full-scale fine-tuning, and red denotes fine-tuning with the ‘Top 5%’ of the model parameters frozen. Dashed lines indicate a learning rate ($lr$) of 5e-6, and solid lines represent $lr$ of 5e-5. We find fine-tuning with the ‘Top 5%’ region frozen during further pre-training effectively mitigates forgetting of non-target languages while maintaining target language acquisition.
4 Realated Work
Neuron Importance Estimating
Several works have empirically estimated neuron importance. An effective importance metric is to utilize parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021). However, such a simple approach may inadequately measure a weight’s impact on model output. Another importance metric is to estimate sensitivity of parameters, essentially approximates the change in loss when a parameter is zeroed out (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023). These studies all focus on pruning unimportant parameters through parameter importance estimation. In this work, we select the most crucial parameters to unveil the core linguistic and monolingual regions.
Cross-lingual Transfer
Multilingual language models exhibit significant zero-shot and few-shot cross-lingual transferability across diverse tasks (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023). Fine-tuned on one language enables model to obtain comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), often displaying code-switching behavior in context generation (Khanuja et al., 2020; Zhao et al., 2024). While enhancements in cross-lingual generalization through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), compulsory language alignment (Sherborne and Lapata, 2022; Shaham et al., 2024) and in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023) have been effective, comprehensive understanding of the internal mechanisms enabling cross-linguistic alignment in large language models (LLMs) is still lacking.
Linguistic Abilities Probing
Researchers have investigated the mechanisms underlying strong cross-lingual performance. Prior works have shown that multilingual multilingual language models rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019). However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that mBERT (Devlin et al., 2019) embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022). Analysis from a hierarchical perspective reveals that classifiers linked to different BERT (Devlin et al., 2019) layers assess semantic features through varied probe tasks (Lin et al., 2019; Jawahar et al., 2019). In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
5 Conclusion
This paper explores the pivotal role of certain parameters in Large Language Models (LLMs), identifying a core region essential for multilingual alignment and generalization. Removing this region causes a complete loss of linguistic ability in LLMs. Further more, we discover that this core region is concentrated in specific dimensions, perturbing only one dimension can cause a significant decrease in language ability. Moreover, beyond the core linguistic regions, we observe that monolingual regions exist within LLMs that affect specific languages. Importantly, we note that the catastrophic forgetting phenomenon during further pre-training may be related to drastic changes in core linguistic regions, as freezing this part during further pre-training alleviates the issue substantially. Our analysis and findings provide new perspectives and explanations for LLMs’ linguistic competence.
Limitations
In this paper, while we discover the core linguistic region and distinct monolingual regions within Large Language Models (LLMs), our work presents two notable limitations. First, our experiments are based on LLaMA-2-7B/13B, and it remains to be further determined whether the same phenomenon are observable in larger or differently architected models. Despite this, our focus on LLaMA-2-7B/13B reveals the existence of linguistic regions within the model, providing an explanation for the model’s linguistic capabilities. Secondly, we optimize full-scale fine-tuning through the freezing operation, which is not suited to extensive datasets. A more feasible approach is to limit the magnitude of parameter updates, which is the direction of our future experiments. Nevertheless, it is important to emphasize that slowing down forgetting through freezing core region suggests that in further pre-training, the core region is different from the other regions. Range of variation amplitude in core region should be smaller to maintain the cross-lingual generalization capabilities of the model. Additionally, while our study focuses on linguistic regions, beyond language, knowledge is a higher-level semantic representation, which is a critical direction for us to explore in the future.
Appendix A Languages in Evaluation Corpus
We use evaluation data composed of 30 languages to assess the model’s linguistic competence. The 30 languages and their respective token counts (use LLaMA-2 Tokenizer) are as follows: Arabic (4702998), Chinese (2869208), Czech (1362041), Danish (36467), Dutch (3991305), English (1216599), Finnish (372303), French (6755281), German (2884921), Greek (474622), Hungarian (1229433), Indonesian (19226), Italian (6332560), Japanese (501899), Korean (2730794), Malay (5842), Malayalam (1489244), Norwegian (42289), Persian (1736589), Polish (4948702), Portuguese (7598161), Romanian (1381598), Russian (5205716), Spanish (7163860), Swahili (630), Swedish (1450236), Tamil (2920808), Turkish (2484186), Ukrainian (455720), Vietnamese (3606202).
Appendix B Evaluation on 30 Languages
The regions are localized from six languages: Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese.
Removal 3% ratio (100K)
LLaMA-2 perplexity on 30 languages when the removal ratio is 3% ratio, with $100,000$ samples for each language. Refer to Table 14 for more details.
Removal 3% ratio (10K)
LLaMA-2 perplexity on 30 languages when the removal ratio is 3% ratio, with reduced $10,000$ samples for each language. Refer to Table 15 for more details.
Removal 1% and 5% ratio (100K)
LLaMA-2-7B perplexity on 30 languages when the removal ratio is changed to 1% and 5% ratio, with $100,000$ equivalent samples for each language. Refer to Table 16 for more details.
Appendix C Remove and Freeze Fine-tuning
In Table 6, we expand removing-freezing ratio to $5\%$ for LLaMA-2-7B fine-tuning, and also observe that removing the ‘Top 5%’ regions destroys the original cross-lingual generalization competence.
Testing Dataset (Language) # Training Samples (Chinese) Removal Ratio $=5\%$
Top & Freeze Bottom & Freeze Top & Unfreeze
Wechat (Chinese) 0K 204630.063 6.454 204630.063
2K 816.94 6.057 6.05
5K 302.806 6.039 6.054
10K 167.891 6.285 6.315
20K 57.355 6.525 6.56
50K 18.92 6.154 6.173
200K 8.8 5.582 5.584
Falcon (English) 0K 66877.711 14.027 66877.711
2K 248218.656 14.486 14.44
5K 476895.25 14.831 15.074
10K 467330.656 16.248 16.334
20K 131648.75 18.579 18.743
50K 25494.166 29.576 29.377
200K 6948.663 321.353 504.645
Table 6: Removing-freezing analysis at $5\%$ removal ratio in different regions of LLaMA-2-7B. ‘Top/Bottom’ denotes the removal region, while ‘Freeze/Unfreeze’ indicates whether the corresponding region is frozen after removal.
Appendix D Attn. Dimensional Removal Evaluation
Figure 6 (left) illustrates that the columns of the Attn.k/q/v matrices in the attention layer, as well as the rows of the Attn.o matrix, correspond to different attention head parameters. Conversely, the rows of the Attn.k/q/v matrices and the columns of the Attn.o matrix are closely associated with features in the representation space.
We remove the ‘Top’ dimensions in the attention layer, and the results is displayed in Tables 7 and 8. Table 7 reveals that removing the ‘Top’ dimensions continues to produce more detrimental effects than other dimensions. The visualizations in Figure 2 show that these dimensions are largely concentrated in a few attention heads, suggesting that some attention heads contribute more significantly to the model’s linguistic competence. Table 8 indicates that the removals under the second setting cause more damage than the first. Considering that, in the second setting, the ‘Top’ dimensions in the matrix directly interact with the corresponding features in the representational space, we can conjecture that these features are tightly linked with the model’s linguistic competence.
Figure 6: One can see from the left that each row of the Attn.o ($W_{o}$) corresponds to a particular attention head, and each column of the Attn.q/k/v ($W_{q/k/v}$) matrix corresponds to one as well. On the right, one can observe the perturbation applied to one weight within RMSNorm, which can be seen as affecting a column of the FFN.down and the Attn.o.
Model Size # Training Samples $N_{d}$ Attn.o(row), Attn.k/q/v(column)
Top Middle Bottom Random
7B 100K $1$ 9.731 6.448 6.445 6.471
100K $3$ 25.82 6.449 6.445 6.474
100K $5$ 62.794 6.452 6.446 6.482
100K $10$ 875.016 6.456 6.446 6.504
13B 100K $1$ 10.899 5.857 5.856 5.856
100K $3$ 44.384 5.858 5.855 5.98
100K $5$ 33.52 5.861 5.856 5.884
100K $10$ 118.968 5.863 5.857 5.966
13B 10K $1$ 8.094 5.856 5.855 5.864
10K $3$ 21.561 5.857 5.855 5.866
10K $5$ 111.766 5.858 5.856 5.865
10K $10$ 108.133 5.861 5.857 5.977
Table 7: Perplexity of LLaMA-2 after removing certain dimensions (zeroed-out) in the attention (Attn) layers. Here, $N_{d}$ denotes the number of dimensions to remove, ’Top’, ’Middle’, and ’Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training across six languages, respectively. ’Random’ denotes an equivalent number of dimensions chosen at random for comparison.
Model Size # Training Samples $N_{d}$ Attn.o(column), Attn.k/q/v(row)
Top Middle Bottom Random
7B 100K $1$ 167.804 6.446 6.446 6.446
100K $3$ 68554.102 6.446 6.447 6.448
100K $5$ 4259.861 6.449 6.447 6.449
100K $10$ 68170.25 6.454 6.452 6.449
13B 100K $1$ 17.609 5.855 5.856 5.856
100K $3$ 313.178 5.857 5.856 5.863
100K $5$ 526.464 5.858 5.856 5.857
100K $10$ 5841.446 5.859 5.858 5.852
13B 10K $1$ 17.03 5.855 5.856 5.857
10K $3$ 206.225 5.856 5.856 5.858
10K $5$ 1110.781 5.857 5.856 5.86
10K $10$ 9600.097 5.859 5.858 5.874
Table 8: Perplexity of LLaMA-2 after removing certain dimensions in attention (Attn) layers. Different from Table 7, in this table, the columns of the Attn.O and the rows of the Attn.K/Q/V are removed.
Appendix E Single Parameter Perturbation
In a Transformer block, each column in the Attn.o and the MLP.down matrix of the feed-forward layer can be considered as the input weights of a neuron. Thus, perturbing a column can be seen as disturbing the input weights of a neuron. Viewed from another angle, if we disturb the output activation value of this neuron, a similar effect should be observed. Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight. Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons (In Figure 6 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix).
Appendix F Individual Language Family Region
Tables 9-13 demonstrate LLaMA-2-7B perplexity after removing Arabic, Spanish, Chinese, Korean, and Vietnamese regions, respectively. The region is obtained by removing the intersections with other languages’ respective regions from the 1% ‘Top/Bottom’ regions, selected from 10,000 or 100,000 sentences during further pre-training according to Equation 4.
Languages Arabic (10K) Arabic (100K)
Base Top Bottom Top Bottom
Arabic 6.771 81.659 6.785 135.02 6.786
Chinese 8.562 9.309 8.593 9.165 8.588
Italian 14.859 16.61 14.959 16.366 14.919
Japanese 10.888 12.238 10.932 11.956 10.923
Korean 4.965 5.534 4.972 5.442 4.969
Persian 6.509 34.142 6.52 43.414 6.508
Portuguese 15.318 16.909 15.262 16.86 15.239
Russian 12.062 13.708 12.145 13.781 12.141
Spanish 17.079 18.543 17.24 18.314 17.2
Ukrainian 9.409 11.243 9.433 11.225 9.439
Vietnamese 5.824 6.412 5.874 6.335 5.871
Table 9: LLaMA-2-7B perplexity on 11 languages with an Arabic region removal. Here, ‘Arabic’ and ‘Persian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged. Values with greater changes compared to the other regions’ removals are in bold.
Languages Spanish (10K) Spanish (100K)
Base Top Bottom Top Bottom
Arabic 6.771 7.158 6.788 7.15 6.789
Chinese 8.562 8.984 8.594 8.971 8.596
Italian 14.859 21.292 14.933 27.004 14.95
Japanese 10.888 11.376 10.913 11.426 10.933
Korean 4.965 5.169 4.967 5.167 4.972
Persian 6.509 6.906 6.484 6.945 6.529
Portuguese 15.318 21.217 15.249 26.877 15.256
Russian 12.062 13.039 12.133 13.252 12.141
Spanish 17.079 38.876 17.224 64.513 17.225
Ukrainian 9.409 10.027 9.439 10.082 9.439
Vietnamese 5.824 6.136 5.875 6.145 5.877
Table 10: LLaMA-2-7B perplexity on 11 languages with a Spanish region removal. Here, ‘Spanish’, ‘Italian’ and ‘Portuguese’ are gray-filled while others are unfilled, and values with greater changes compared to the other regions’ removals are in bold.
Languages Chinese (10K) Chinese (100K)
Base Top Bottom Top Bottom
Arabic 6.771 7.161 6.79 7.714 6.784
Chinese 8.562 10.899 8.592 12.079 8.586
Italian 14.859 16.041 14.939 15.881 14.932
Japanese 10.888 12.265 10.922 12.878 10.904
Korean 4.965 5.343 4.974 5.341 4.960
Persian 6.509 6.92 6.519 6.865 6.516
Portuguese 15.318 16.285 15.27 16.241 15.26
Russian 12.062 12.887 12.136 12.973 12.145
Spanish 17.079 18.068 17.216 17.974 17.219
Ukrainian 9.409 10.144 9.439 10.207 9.447
Vietnamese 5.824 6.261 5.878 6.296 5.870
Table 11: LLaMA-2-7B perplexity on 11 languages with a Chinese region removal. Here, ‘Chinese’ and ‘Japanese’ are gray-filled while others are unfilled, and values with greater changes compared to the other regions’ removals are in bold.
Languages Korean (10K) Korean (100K)
Base Top Bottom Top Bottom
Arabic 6.771 7.259 6.791 7.316 6.783
Chinese 8.562 9.14 8.594 9.173 8.594
Italian 14.859 15.91 14.941 15.791 14.938
Japanese 10.888 13.273 10.919 15.062 10.932
Korean 4.965 8.364 4.971 13.128 4.971
Persian 6.509 7.38 6.522 7.574 6.522
Portuguese 15.318 16.113 15.259 15.984 15.26
Russian 12.062 12.758 12.138 12.827 12.136
Spanish 17.079 17.981 17.214 17.858 17.225
Ukrainian 9.409 10.065 9.434 10.108 9.442
Vietnamese 5.824 6.188 5.874 6.177 5.874
Table 12: LLaMA-2-7B perplexity on 11 languages with a Korean region removal. Here, ‘Korean’ and ‘Japanese’ are gray-filled while others are unfilled, and values with greater changes compared to the other regions’ removals are in bold.
Languages Vietnamese (10K) Vietnamese (100K)
Base Top Bottom Top Bottom
Arabic 6.771 7.435 6.785 7.341 6.789
Chinese 8.562 9.576 8.589 9.372 8.592
Italian 14.859 16.979 14.952 16.497 14.937
Japanese 10.888 12.027 10.946 11.814 10.941
Korean 4.965 5.44 4.97 5.335 4.979
Persian 6.509 7.315 6.501 7.243 6.521
Portuguese 15.318 17.159 15.249 16.805 15.258
Russian 12.062 13.107 12.141 13.007 12.144
Spanish 17.079 18.801 17.244 18.369 17.233
Ukrainian 9.409 10.316 9.447 10.217 9.433
Vietnamese 5.824 24.382 5.872 27.817 5.874
Table 13: LLaMA-2-7B perplexity on 11 languages with a Vietnamese region removal. Here, ‘Vietnamese’ is gray-filled while others are unfilled, and values with greater changes compared to the other regions’ removals are in bold.
Languages LLaMA-2-7B 3% (100K) LLaMA-2-13B 3% (100K)
Base Top Bottom Random Base Top Bottom Random
Arabic 6.771 127208.250 6.772 7.895 6.261 102254.758 6.316 7.112
Chinese 8.652 295355.5 8.565 9.837 7.838 84086.906 7.806 8.619
Czech 19.834 62692.367 19.835 24.005 17.744 56102.227 17.650 20.485
Danish 8.372 47654.156 8.372 9.929 7.402 47213.586 7.401 8.278
Dutch 16.959 48478.594 16.959 20.121 15.64 46303.559 15.572 18.295
English 7.653 16573.422 7.653 8.359 7.447 25212.217 7.234 7.821
Finnish 7.566 45711.992 7.566 8.934 6.887 48811.242 6.861 7.826
French 13.605 48268.211 13.605 15.003 12.765 45674.492 12.573 13.682
German 18.355 64015.117 18.356 15.404 17.29 51692.125 16.973 18.972
Greek 3.832 224595.781 3.833 4.527 3.599 80657.891 3.599 4.146
Hungarian 16.365 52828.691 16.363 20.039 14.756 58107.137 14.834 17.633
Indonesian 44.269 33121.945 44.318 48.175 37.909 51611.625 37.838 38.548
Italian 14.859 58908.879 14.860 17.341 13.694 47375.844 13.730 15.207
Japanese 10.888 322031.406 10.896 12.535 10.072 75236.031 10.137 11.661
Korean 4.965 125345.359 4.967 5.649 4.724 90768.844 4.743 5.241
Malay 66.581 22603.727 66.843 74.167 46.885 40468.750 46.912 58.947
Malayalam 5.133 373710.188 5.134 6.396 4.972 16990.266 4.972 5.654
Norwegian 14.425 31526.176 14.427 17.854 13.142 45820.109 13.139 15.041
Persian 6.509 81959.719 6.511 7.628 6.205 92201.812 6.229 7.009
Polish 12.629 66906.469 12.629 14.843 11.414 55923.156 11.311 12.987
Portuguese 15.318 47763.059 15.319 17.297 13.667 51498.402 13.982 15.376
Romanian 10.893 43498.008 10.895 13.061 9.652 54986.055 9.693 10.969
Russian 12.062 170776.750 12.064 13.728 11.048 112574.609 10.948 11.757
Spanish 17.079 51940.859 17.082 18.98 16.351 54005.891 16.138 17.292
Swahili 75.908 29234.168 75.892 89.380 70.519 48802.227 70.402 81.216
Swedish 14.714 49425.969 14.714 17.258 13.229 48622.266 13.337 14.933
Tamil 4.162 381070.844 4.162 5.04 4.028 111060.516 4.049 4.488
Turkish 11.214 46986.391 11.215 13.765 9.834 50303.562 9.763 11.374
Ukrainian 9.409 120719.938 9.409 10.875 8.295 116287.305 8.297 9.076
Vietnamese 5.824 40126.527 5.824 6.614 5.471 42336.426 5.437 5.995
Table 14: LLaMA-2 perplexity on 30 languages with 3% removal ratio. ‘100K’ means that the region is selected from 100,000 samples. ‘Top’ and ‘Bottom’ respectively indicate the $N$ parameters with the highest and lowest cumulative $\mathcal{I}^{*}_{j}(\theta)$ during the further pre-training across the six languages. ‘Random’ denotes the randomly selecting $N$ while ‘Base’ represents no removal. Here, $N$ equals 3% of the total number in each parameter matrix.
Languages LLaMA-2-7B 3% (10K) LLaMA-2-13B 3% (10K)
Base Top Bottom Random Base Top Bottom Random
Arabic 6.771 115398.328 6.772 7.895 6.261 88678.016 6.315 7.112
Chinese 8.652 369027.531 8.564 9.837 7.838 70912.242 7.806 8.619
Czech 19.834 78480.219 19.837 24.005 17.744 53699.43 17.655 20.485
Danish 8.372 46503.742 8.373 9.929 7.402 39408.91 7.401 8.278
Dutch 16.959 55704.191 16.961 20.121 15.64 41159.938 15.572 18.295
English 7.653 15738.043 7.654 8.359 7.447 23678.322 7.234 7.821
Finnish 7.566 46616.094 7.568 8.934 6.887 47002.539 6.861 7.826
French 13.605 44385.668 13.609 15.003 12.765 38755.539 12.678 13.682
German 18.355 84497.234 18.361 21.404 17.29 43319.586 17.02 18.972
Greek 3.832 147740.5 3.833 4.527 3.599 70136.242 3.6 4.146
Hungarian 16.365 52652.363 16.367 20.039 14.756 48407.305 14.735 17.633
Indonesian 44.269 39055.945 44.267 48.175 37.909 36912.34 37.929 38.548
Italian 14.859 54297.523 14.865 17.341 13.694 42515.969 13.69 15.207
Japanese 10.888 358722.188 10.891 12.535 10.072 68055.984 10.118 11.661
Korean 4.965 102918.828 4.966 5.649 4.724 65209.328 4.736 5.241
Malay 66.581 23501.082 67.158 74.167 46.885 35517.879 47.191 58.947
Malayalam 5.133 314088.969 5.136 6.396 4.972 131629.438 4.971 5.654
Norwegian 14.425 38111.27 14.431 17.854 13.142 38500.664 13.138 15.041
Persian 6.509 78203.031 6.51 7.628 6.205 98292.281 6.22 7.009
Polish 12.629 81373.273 12.633 14.843 11.414 52403.461 11.393 12.987
Portuguese 15.318 47779.789 15.321 17.297 13.667 41184.457 13.86 15.376
Romanian 10.893 45836.578 10.897 13.061 9.652 51766.957 9.694 10.969
Russian 12.062 227916.828 12.061 13.728 11.048 103490.719 11.004 11.757
Spanish 17.079 57679.461 17.087 18.98 16.351 40338.426 16.265 17.292
Swahili 75.908 42977.977 75.93 89.38 70.519 40400.949 70.443 81.216
Swedish 14.714 55893.812 14.717 17.258 13.229 45396.66 13.301 14.933
Tamil 4.162 447989.969 4.162 5.04 4.028 141214.188 4.052 4.488
Turkish 11.214 57037.605 11.215 13.765 9.834 41566.105 9.791 11.374
Ukrainian 9.409 168085.672 9.408 10.875 8.295 94307.312 8.296 9.076
Vietnamese 5.824 36374.734 5.825 6.614 5.471 31730.328 5.467 5.995
Table 15: LLaMA-2 perplexity on 30 languages with 3% removal ratio. ‘10K’ means that the region is selected from 10,000 samples. Here, we reduce training samples to $10,000$ during further pre-training across six languages.
Languages LLaMA-2-7B 1% (100K) LLaMA-2-7B 5% (100K)
Base Top Bottom Random Base Top Bottom Random
Arabic 6.771 67579496 6.77 7.021 6.771 112504.609 6.774 10.823
Chinese 8.652 120887480 8.561 8.818 8.652 156026.938 8.565 12.775
Czech 19.834 24343856 19.835 21.176 19.834 96580.281 19.845 24.797
Danish 8.372 1631186.625 8.372 8.775 8.372 82876.266 8.375 13.565
Dutch 16.959 6845146 16.963 18.056 16.959 79497.211 16.961 27.01
English 7.653 512756 7.654 7.851 7.653 46197.477 7.656 9.289
Finnish 7.566 4727027.5 7.567 7.948 7.566 60183.328 7.56 13.005
French 13.605 4768049 13.608 14.198 13.605 87642.109 13.611 19.076
German 18.355 17940508 18.357 19.724 18.355 106160.992 18.364 28.772
Greek 3.832 14242545 3.833 3.972 3.832 141320.578 3.835 6.45
Hungarian 16.365 130584 16.366 17.35 16.365 77265.188 16.369 30.376
Indonesian 44.269 1654245 44.347 49.476 44.269 83353.344 44.298 64.743
Italian 14.859 5265871.5 14.863 15.607 14.859 83076.164 14.865 22.6
Japanese 10.888 28104000 10.88 11.196 10.888 124647.633 10.895 16.619
Korean 4.965 16449047 4.965 5.095 4.965 59954.559 4.967 7.831
Malay 66.581 7875206 66.673 78.545 66.581 51824.859 66.751 90.933
Malayalam 5.133 7151096 5.133 5.359 5.133 182008.484 5.137 7.905
Norwegian 14.425 4223085 14.429 15.35 14.425 79399.109 14.434 23.621
Persian 6.509 2233196 6.507 6.782 6.509 107342.734 6.511 10.236
Polish 12.629 6547834.5 12.631 13.36 12.629 88912.945 12.632 21.372
Portuguese 15.318 6249820 15.319 15.927 15.318 78851.766 15.324 22.608
Romanian 10.893 5251915.5 10.895 11.526 10.893 71228.375 10.899 19.21
Russian 12.062 17596800 12.061 12.067 12.062 102639.602 12.066 18.504
Spanish 17.079 8220832.5 17.084 18.029 17.079 96575.547 17.084 24.007
Swahili 75.908 7875009 75.845 83.963 75.908 77765.133 75.8 131.709
Swedish 14.714 4712167.5 14.716 15.534 14.714 81574.734 14.717 22.628
Tamil 4.162 20660974 4.162 4.265 4.162 173728.312 4.164 5.881
Turkish 11.214 4489915 11.214 11.882 11.214 58347.055 11.218 19.76
Ukrainian 9.409 11689088 9.409 9.811 9.409 90008.312 9.414 14.807
Vietnamese 5.824 2235468 5.825 6.018 5.824 54187.02 5.825 9.015
Table 16: LLaMA-2-7B perplexity on 30 languages with 1% and 5% removal ratio. ‘100K’ means that the region is selected from 100,000 samples. Here, we change the removal ratio from 3% to 1% and 5%.

### v2

@@
 Furthermore, this core region exhibits significant dimensional dependency, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence.
-Moreover, we discover that distinct regions exist for different monolingual families, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages.
-Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common occurrence observed during further pre-training of LLMs.
+Moreover, we discover that distinct monolingual regions exist for different languages, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages.
+Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common phenomenon observed during further pre-training of LLMs.
 Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence.
 Unveiling Linguistic Regions in Large Language Models
-[…]
+Zhihao Zhang1thanks:  Equal contributions., Jun Zhao1, Qi Zhang13thanks:  Corresponding authors., Tao Gui2, Xuanjing Huang1 1 School of Computer Science, Fudan University 2 Institute of Modern Languages and Linguistics, Fudan University 3 Shanghai Collaborative Innovation Center of Intelligent Visual Computing {zhangzhihao19, zhaoj19, qz, tgui, xjhuang}@fudan.edu.cn
 1 Introduction
@@
 Over the years, the field of Natural Language Processing (NLP) has been at the forefront of understanding the core principles of intelligence.
-The emergence of Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023), and LLaMA 2 (Touvron et al., 2023), showcases a significant breakthrough.
+The emergence of Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023) and LLaMA 2 (Touvron et al., 2023), showcase a significant breakthrough.
 Thanks to unparalleled scales of model architecture and the vastness of training data, these LLMs now exhibit exceptional linguistic competence and can execute complex tasks requiring abstract knowledge (Dong et al., 2023) and reasoning (Cobbe et al., 2021).
 Previous research has revealed that LLMs naturally capture cross-linguistic similarities in their representation space, facilitating zero-shot cross-lingual transfer (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023).
-The model is fine-tuned on one language, enabling the acquisition of comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), and exhibits the phenomenon of code-switching when generating context, switching between languages within a single utterance (Khanuja et al., 2020; Zhao et al., 2024).
+The model is fine-tuned on one language, enabling the acquisition of comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), and exhibits the phenomenon of code-switching when generating context (Khanuja et al., 2020; Zhao et al., 2024).
 Attempts to improve LLMs’ cross-lingual generalization abilities have been successful through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), aligning languages compulsorily (Sherborne and Lapata, 2022; Shaham et al., 2024) and utilizing in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023).
-However, a detailed research of underlying the internal mechanisms of how LLMs possess cross-linguistic alignment capability remains elusive.
-To delve deeper into the intrinsic mechanisms of LLMs’ linguistic competence, this paper focus on the LLMs’ parameter importance and investigate the linguistic regions of LLMs based on 30 distinct languages’ performance, with the purpose of figuring out the following questions:
+However, a detailed investigation into the internal mechanisms of how LLMs possess cross-linguistic alignment capability remains elusive.
+To delve deeper into the intrinsic mechanisms of LLMs’ linguistic competence, this paper focuses on the LLMs’ parameter importance and investigate the linguistic regions of LLMs based on 30 distinct languages’ performance, with the purpose of figuring out the following questions:
 Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual generalization?
-Through pre-training across six languages and measurement of parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters.
-As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently lead to a significant decline in performance across 30 test languages (Section 3.2).
+By conducting further pre-training across six languages and evaluating models’ parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters.
+As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently leads to a significant decline in performance across 30 test languages (Section 3.2).
 Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependency.
 In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3).
-Q2: Beyond the core multilingual region within LLMs , do distinct monolingual (or monolingual family) regions exist that specifically influence individual languages?
-While LLMs possess strong multilingual capabilities, we discover that each individual language (or language family) encompasses independent regions within the LLMs.
+Additionally, ablation study in 3.3 shows that beyond outlier dimensions, other non-outlier dimensions in this region are also critical.
+Q2: Beyond the core linguistic region within LLMs , do distinct monolingual regions exist that specifically influence individual languages?
+While LLMs possess strong multilingual capabilities, we discover that each individual language (or language with similar compositional elements or grammatical structures) encompasses independent regions within the LLMs.
 As shown in the middle of Figure 1, the analysis of the Russian sentences identifies a particular linguistic region that likewise exerts influence both on the Russian and Ukrainian language, both of which belong to the Slavic group (Section 3.4).
@@
 Notable shifts in these regions potentially lead to a decline in model lingual capabilities.
-Our findings reveal that freezing this core region can mitigate the issue of catastrophic forgetting (McCloskey and Cohen, 1989; Kemker et al., 2018), a common occurrence observed during further pre-training of LLMs.
-As shown at the bottom of Figure 1, we investigate the impact of selectively freezing 5% key parameters of all parameters fine-tuning during further pre-training, compared to the full-scale fine-tuning technique.
-Findings indicate that this method facilitated comparable or even more efficient learning of the target language, while concurrently decelerating the rate of language attrition for previously learned languages (Section 5).
-Significantly, our methodology is compatible with the data-replay techniques (Robins, 1995), with no necessity for integrating extra components into the model.
+Our findings reveal that freezing this core region can mitigate the issue of catastrophic forgetting (McCloskey and Cohen, 1989; Kemker et al., 2018), a common phenomenon observed during further pre-training of LLMs.
+As shown at the bottom of Figure 1, we investigate the impact of selectively freezing 5% key parameters of all parameters during further pre-training, compared to the full-scale fine-tuning technique.
+Findings indicate that this method facilitates comparable learning of the target language while concurrently decelerating the rate of language attrition for previously learned languages (Section 3.5).
+Significantly, our methodology is compatible with the data-replay techniques (Robins, 1995; Wei et al., 2023), with no necessity for integrating extra components into the model.
 Unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), our approach restricts to a minimal core region in LLMs.
@@
 Furthermore, perturbations to specific dimensions or even a single parameter can lead to a substantial decline in the model’s linguistic abilities.
-  • We observe that distinct monolingual regions exist in LLMs for different languages (or language families).
-Removing a specific monolingual region causes a significant deterioration in the linguistic capabilities within that language (or language family).
-  • We perform further pre-training for specific languages within the core linguistic region of LLMs frozen, achieving comparable or even superior performance in the target language while mitigating catastrophic forgetting in non-target languages.
+  • We observe that distinct monolingual regions exist in LLMs for different languages.
+Removing a specific monolingual region causes a significant deterioration in the linguistic capabilities within corresponding language.
+  • We perform further pre-training for specific languages within the core linguistic region of LLMs frozen, achieving comparable performance in the target language while mitigating catastrophic forgetting in non-target languages.
 2 Background and Metric
@@
 Pre-training is a crucial process by which LLMs acquire linguistic competence and gain general knowledge about the real world.
-Formally, given a large corpus $\mathcal{D}$, the training objective is to find the optimal $\theta$, minimize the following loss $\mathcal{L}$:
+Formally, given a large corpus $\mathcal{D}$, the training objective for auto-regressive language modeling is to find the optimal $\theta$ that minimizes the following loss $\mathcal{L}$:
 $\mathcal{L}{(\mathcal{D},\theta)}=\sum_{x\in\mathcal{D}}\sum_{i}\log p_{\theta}(x_{i}|x_{1},...,x_{i-1}),$ (1)
@@
 This becomes particularly challenging as the number of model parameters, $d$, grows to hundreds of billions.
-However, according to the Taylor expansion formula for $\mathcal{L}$ at $\theta_{j}=0$:
+However, similar to several prior works (Molchanov et al., 2019; Zhang et al., 2022), using the Taylor expansion formula for $\mathcal{L}$ at $\theta_{j}=0$:
 $\quad\mathcal{L}(\mathcal{D},\theta)=\mathcal{L}(\mathcal{D},\theta|\theta_{j}=0)$ (3)
@@
 3.2 Core Linguistic Competence Region
-In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify core parameter regions associated with linguistic competence.
-Specifically, according to Equation 4, we cumulatively compute *$\mathcal{I}^{*}(\theta)=\Sigma\mathcal{I}(\theta)$ values across six different languages’ training, positing that the set of parameters exhibiting maximal importance score *$\mathcal{I}^{*}(\theta)$ during the language further pre-training may have a strong correlation with the model’s linguistic competence, and we provide both logical and empirical evidence to support this hypothesis.
+In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify the core linguistic region in LLMs.
+Here, we define it as "core linguistic", attributed to its minimal proportion within LLMs, constituting only 1% of the total parameters, and its association with model’s linguistic competence modeling.
+Specifically, according to Equation 4, we cumulatively compute $\mathcal{I}^{*}(\theta)=\Sigma\mathcal{I}(\theta)$ values across six different languages’ training, positing that the set of parameters exhibiting maximal importance score $\mathcal{I}^{*}(\theta)$ during the language further pre-training may have a strong correlation with the model’s linguistic competence.
+We provide both logical and empirical evidence to support this hypothesis.
 Logical Evidence
@@
 Figure 2: Visualization of the linguistic competence region (the ‘Top’ 5% region).
-The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the Top region.
+The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the ‘Top’ region.
 Testing Dataset (Language) # Training Samples (Chinese) Removal Ratio $=1\%$
@@
 This implies that the ‘Top’ region encodes generalizable fundamental linguistic competence.
-When ‘Top’ region is zeroed-out and frozen, other regions significantly adapt to regain core linguistic competence.
-Similar conclusions can be obtained with an expanded removal ratio 5%.
-For further details, see Appendix C.
+When ‘Top’ region is zeroed-out and frozen, other regions difficultly adapt to regain core linguistic competence.
 Model Size # Training Samples $N_{d}$ Attn.o(row), Attn.k/q/v+FFN.down(column)
@@
 Table 3: Perplexity of LLaMA-2 after removing certain dimensions in the Attention and Feedforward layers.
-Here, $N_{d}$ denotes the number of dimensions to remove, ’Top’, ’Middle’, and ’Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training.
-’Random’ denotes an equivalent number of dimensions chosen randomly for comparison.
+Here, $N_{d}$ denotes the number of dimensions to remove, ‘Top’, ‘Middle’, and ‘Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training.
+‘Random’ denotes an equivalent number of dimensions chosen randomly for comparison.
 3.3 Dimensional Dependence
@@
 As shown in Figure 2, whether in the attention mechanism layer or the feed-forward layer, the linguistic region displays a distinct concentration in both the rows and columns of the matrices.
-Such distribution features seem to imply that the model’s linguistic competence is concentrated in specific dimensions.
+In Appendix B, we also discover that in various layers, the core linguistic region is concentrated on different heads of the Attn.o matrix.
+Such distribution features seem to imply that the model’s linguistic competence is localized in specific rows and columns.
 Structured Removal
@@
 However, disturbances to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects.
-Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in D.
+Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in C.
 Single Dimension Perturbation
-Here, we explore whether a specific dimension significantly impact the model’s linguistic competence.
-As illustrated in Figure 3, we iterate through the key dimensions mentioned in Section 3.3, attempting to perturb the same dimension (random initialization) across all Transformer layers.
-The results indicate that the impact of dimensions 2100 and 4743 on the LLaMA-2-13B substantially surpassed other dimensions, even when compared to the other three in the ‘Top5’ dimensions.
-In contrast, perturbing two randomly selected dimensions, 2800 and 4200, yield linguistic performance almost indistinguishable from the unperturbed state.
-Figure 3: Perplexity of the LLaMA-2-13B when perturbing the same single dimension (Att.O and FFN.down matrices) across all layers.
+Here, we explore whether a specific dimension significantly impacts the model’s linguistic competence.
+As illustrated in Figure 3, we iterate through the key dimensions mentioned in Section 3.3, attempting to perturb (random initialization) the same dimension across all Transformer layers.
+The results indicate that the impact of dimensions 2100 and 4743 on the LLaMA-2-13B substantially surpasses other dimensions, even when compared to the other three in the ‘Top5’ dimensions.
+In contrast, perturbing two randomly selected dimensions, 2800 and 4200, yields linguistic performance almost indistinguishable from the unperturbed state.
+Figure 3: Perplexity of the LLaMA-2-13B when perturbing the same single dimension (Att.O and FFN.down columns) across all layers.
 ‘Top k’ represents the top k dimensions that disrupt the model the most.
@@
 If this weight parameter is multiplied by $10$, the PPL value also rises to $4363.462$.
-This suggests that even minor changes to a single parameter can cause the model to lose nearly all of its linguistic competence.
 However, randomly altering the parameters at dimensions 2800 and 4200 doesn’t noticeably impact the model.
-For more details, refer to Appendix E.
+For more details, refer to Appendix D.
 Output Under Perturbations
@@
 The results are shown in Figure 4.
-Compared to randomly selected dimension 4200-th, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
-3.4 Monolingual Family Region
-In this section, we wonder if LLMs possess distinct regions within an individual language (or language family).
+Compared to randomly selected 4200-th dimension, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
+Ablation study
+Considering that the collapse of PPL may be possibly caused solely by the removal of outlier dimension parameters, rather than the collective effect of the entire linguistic region, we conduct an ablation experiment without removing the outlier dimensions’ parameters (1512/2533 for LLaMA-2-7B and 2100/4743 for LLaMA-2-13B).
+The results for the 13B model, presented in Table 4, indicate that while outlier dimensions are contained within the core linguistic region, non-outlier dimensions in this region are also critical.
+Not altering any row or column of the matrix parameters for the two outlier dimensions also results in a great increase in PPL, although the collapse of PPL is slowed.
+This finding reveals that all dimensions in the core linguistic region are tightly interrelated, and disrupting even a small part of it can lead to a PPL collapse.
+For a more detailed analysis of the experiments and results for the 7B model, please refer to Appendix E.
+Removal Region $N_{d}$ LLaMA-2-13B Top(100K)
+w/ outlier $d$ w/o outlier $d$
+Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) 1 5218.1 11.079
+3 37344.078 77.519
+5 41840.613 590.895
+10 62579.516 513998.437
+Attn.o(row) Attn.k/q/v(column) 1 10.899 10.901
+3 44.384 44.389
+5 33.52 29.793
+10 118.968 120.977
+Attn.o(column) Attn.k/q/v(row) 1 17.609 13.666
+3 313.178 63.99
+5 526.464 163.388
+10 5841.446 2675.347
+FFN.up/gate(row) FFN.down(column) 1 154.925 6.142
+3 33995.949 6.668
+5 32572.888 8.139
+10 524867.687 45.408
+Table 4: Perplexity of LLaMA-2-13B after removing ‘Top’ certain dimensions w/ or w/o outlier dimensions respectively.
+Here, $N_{d}$ denotes the number of dimensions to remove, ‘Top’ refers to the dimensions with the most cumulated $\mathcal{I}_{\theta}$ during further pre-training.
+Figure 4: Comparison of linguistic competence. Expanding a single parameter to four times leads to error language competence in LLaMA-2-13B, a 13 billion-parameter LLM.
+3.4 Monolingual Region
+In this section, we wonder if LLMs possess distinct regions within an individual language.
 Unlike the core linguistic regions described in Section 3.2, a monolingual family region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.
-Perturbation Parameter Perplexity
-- - 5.865
-Reset 1 L1-N2100 83224.078
-Reset 1 L1-N2800 5.860
-Reset 1 L1-N4200 5.858
-Mul 10 L1-N2100 4363.462
-Mul 10 L1-N2800 5.859
-Mul 10 L1-N4200 5.864
-Table 4: Perplexity of LLaMA-2-13B on Chinese when perturbing a single weight parameter.
-Here, ‘Reset $1$’ represents resetting the parameter to $1$ (the initial value before pre-training), ‘Mul 10’ represents multiplying the parameter by $10$.
-‘L1’ represents $1$-st layers.
-‘N’ represents the ‘Input_LayerNorm’ module, followed by the perturbed dimension.
 Region Localization
@@
 $\mathcal{S}_{l}^{*}=\mathcal{S}_{l}\,\,-\bigcup_{l^{{}^{\prime}}\in L\setminus\{l\}}\mathcal{S}_{l^{{}^{\prime}}}.$ (5)
+In Appendix F, we visualize the distribution Attn.q matrix in ‘Arabic’ and ‘Vietnamese’ regions and discover minimal overlap between them.
 Region Removal
@@
 A similar phenomenon is observed if removals are changed to the regions for each of other five languages, see Appendix F for more details.
-Figure 4: Comparison of linguistic competence. Expanding a single parameter to four times leads to error language competence in LLaMA-2-13B, a 13 billion-parameter LLM.
-3.5 Further Pre-training Optimization
-In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while preserving performance comparable to full-scale fine-tuning in target language proficiency.
-Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of $256$, a maximum token count of $512$, and learning rates ($lr$) and $5e{-5}$ or $5e{-6}$, employing perplexity (PPL) as the evaluation criterion.
-Full-scale Model Fine-tuning
-Traditional full-scale fine-tuning, when increasing the learning rate, enhances learning in the target language but aggravates forgetting in non-target languages.
-To counteract this forgetting, it is often essential to incorporate a portion of data from these other languages.
-As depicted on the left side of Figure 5 in line Blue, increasing $lr$ from $5e{-6}$(dotted blue line) to $5e{-5}$ (solid blue line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while accelerates the forgetting rate of the non-target languages (English and Chinese) simultaneously, shown in the middle and right side.
-Freeze Core Regions Fine-tuning
-We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment.
-To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 5.
-At a $lr$ of $5e{-6}$ (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal.
-However, when the $lr$ increases to $5e{-5}$ (solid line), freezing fine-tuning not only facilitates faster learning in the target language (achieving better performance in Arabic PPL: 3.557 vs.
-3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs.
-20.557 and 90.84 vs.
-563.423 , respectively).
-The potential reason for this phenomenon may lie in the preservation of the core regions within the cross-lingual alignment competence.
-Restricting the magnitude of updates in the core region’s parameters is a future strategy we intend to employ.
-Notably, unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), such approaches restricts to a minimal core region in LLMs, and can be implemented alongside blending previous data, retraining the entire network, or possibly only the final layers, without adding additional components to the model.
 Languages Russian (10K) Russian (100K)
@@
 Table 5: LLaMA-2-7B perplexity on 11 languages with a Russian region removal.
-Here, ‘Arabic’ and ‘Persian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged.
+Here, ‘Russian’ and ‘Ukrainian’ are gray-filled while others are unfilled, ‘Top’ and ‘Bottom’ are deduplicated, and ‘Base’ is unchanged.
 Values with greater changes compared to the other regions’ removals are in bold.
+Downstream Task
+We conduct downstream tasks on MMLU (Hendrycks et al., 2021) and ArabicMMLU (Koto et al., 2024).
+The former is in English while the latter is in Arabic.
+Our experiments demonstrate that removing monolingual regions significantly impacts the model’s ability to perform downstream tasks in the targeted language.
+Specifically, our findings are as follows:
+1) The average logits probability for option ‘A’ on the ArabicMMLU evaluation increases from 0.36 to 0.64.
+2) For the evaluation where the correct answer is one of the options ‘B/C/D/E’ in ArabicMMLU, the model’s accuracy decreases dramatically from 25.6% to just 1.5% when the ‘Arabic’ region is removed.
+However, if the ‘Vietnamese’ region is removed, the model’s accuracy remains almost unaffected at 26.7%.
+3) For the MMLU evaluation, the model’s accuracy is minimally impacted by the removal of monolingual regions.
+Compared to a baseline accuracy of 42.46%, the removals of the ‘Arabic’ and ‘Vietnamese’ regions result in accuracies of 39.27% and 39.68%, respectively.
+Furthermore, using "There are 365 days in a year and 12" as a prompt for generation, we test outputs on the removal of the removing monolingual regions. More details are in Appendix F.
+Figure 5: Model’s accuracy on MMLU and ArabicMMLU (filtered) test. Here, ‘filtered’ denotes removing questions whose correct answer is ‘A’.
+3.5 Further Pre-training Optimization
+In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while maintaining learning proficiency comparable to full-scale fine-tuning in target language.
+Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of $256$, a maximum token count of $512$, and learning rates ($lr$) and $5e{-5}$ or $5e{-6}$, employing perplexity (PPL) as the evaluation criterion.
+Full-scale Model Fine-tuning
+Traditional full-scale fine-tuning, when increasing the learning rate or the amount of corpus data, enhances learning in the target language but aggravates forgetting in non-target languages.
+As shown in Table 2, since LLaMA-2 is primarily trained on English corpora, conducting a second stage of pretraining solely on large-scale Chinese corpora can lead to the forgetting of English competence.
+Additionally, as depicted on the left side of Figure 6 in blue line, increasing $lr$ from $5e{-6}$ (dotted line) to $5e{-5}$ (solid line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while simultaneously accelerates the forgetting rate of the non-target languages (English and Chinese), as shown in the middle and right side.
+Freeze Core Regions Fine-tuning
+We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment.
+To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 6.
+At a $lr$ of $5e{-6}$ (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal.
+However, when the $lr$ increases to $5e{-5}$ (solid line), freezing fine-tuning not only facilitates faster learning in the target language (preserving comparable performance in Arabic PPL: 3.557 vs.
+3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs.
+20.557 and 90.84 vs.
+563.423, respectively).
+The potential reason for this phenomenon may lie in the preservation of the core regions within the cross-lingual alignment competence.
+Restricting the magnitude of updates in the core region’s parameters is a future strategy we intend to employ.
+Notably, unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), such approaches restricts to a minimal core region in LLMs, and can be implemented alongside blending previous data, retraining the entire network, or possibly only the final layers, without adding additional components to the model.
 Figure 6: Perplexity of LLaMA-2 across Arabic, English, and Chinese when training on 100,000 Arabic sentences.
@@
 4 Realated Work
-Neuron Importance Estimating
-Several works have empirically estimated neuron importance.
-An effective importance metric is to utilize parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021).
-However, such a simple approach may inadequately measure a weight’s impact on model output.
-Another importance metric is to estimate sensitivity of parameters, essentially approximates the change in loss when a parameter is zeroed out (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023).
-These studies all focus on pruning unimportant parameters through parameter importance estimation.
-In this work, we select the most crucial parameters to unveil the core linguistic and monolingual regions.
+Intrinsic Regions
+Prior works aimed to extract a sub-network capable of executing specific downstream tasks (Frankle and Carbin, 2019; Zhang et al., 2022) or task-specific subspaces to limit parameter fine-tuning within it (Aghajanyan et al., 2021; Zhang et al., 2023).
+For parameter importance estimation, an effective metric is to use parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021).
+Another metric involves estimating the sensitivity of parameters (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023).
+In this work, we employ the latter method to select the most crucial parameters to unveil linguistic regions.
+(Frankle and Carbin, 2019) extracted a complete sub-network for downstream tasks, as exemplified by the optimal scale of the lottery network at 21.2%.
+However, our findings indicate that the linguistic regions is a much smaller (1%) functional region.
 Cross-lingual Transfer
-Multilingual language models exhibit significant zero-shot and few-shot cross-lingual transferability across diverse tasks (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023).
+Multilingual language models exhibit significant zero-shot and few-shot cross-lingual transferability across diverse tasks (Pires et al., 2019; Xu et al., 2023).
 Fine-tuned on one language enables model to obtain comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), often displaying code-switching behavior in context generation (Khanuja et al., 2020; Zhao et al., 2024).
-While enhancements in cross-lingual generalization through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), compulsory language alignment (Sherborne and Lapata, 2022; Shaham et al., 2024) and in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023) have been effective, comprehensive understanding of the internal mechanisms enabling cross-linguistic alignment in large language models (LLMs) is still lacking.
+While enhancements in cross-lingual generalization through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), compulsory language alignment (Sherborne and Lapata, 2022; Shaham et al., 2024) and in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023) have been effective, a comprehensive understanding of the internal mechanisms enabling cross-linguistic alignment in LLMs is still lacking.
 Linguistic Abilities Probing
-Researchers have investigated the mechanisms underlying strong cross-lingual performance.
-Prior works have shown that multilingual multilingual language models rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019).
+Prior works have shown that multilingual LMs rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019).
 However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that mBERT (Devlin et al., 2019) embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022).
@@
 In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
+Outlier Dimensions
+Multiple studies have identified outlier dimensions in Transformer-based LMs (Kovaleva et al., 2021; Dettmers et al., 2022).
+Researches have found that certain outlier dimensions in PLMs are highly sensitive to the fine-tuning of downstream tasks (Kovaleva et al., 2021; Puccetti et al., 2022).
+Furthermore, (Rudman et al., 2023; Puccetti et al., 2022) discovered that these outlier dimensions encode task-specific knowledge, and disabling these dimensions significantly degrades model performance.
+(Luo et al., 2021) attribute positional embeddings to the emergence of outliers, while (Rajaee and Pilehvar, 2022) reported inconsistent results.
+Additionally, (Sun et al., 2024) found that these outlier dimensions exhibit significantly larger activation values than others in LLMs.
+Our findings in Section 3.3 demonstrate that beyond the outlier dimensions, other non-outlier dimensions within the core linguistic region also play an indispensable role in the model’s core linguistic competence.
 5 Conclusion
@@
 Additionally, while our study focuses on linguistic regions, beyond language, knowledge is a higher-level semantic representation, which is a critical direction for us to explore in the future.
+Acknowledgements
+The authors wish to thank the anonymous reviewers for their helpful comments.
+This work was partially funded by National Natural Science Foundation of China (No.62076069,62206057), Shanghai Rising-Star Program (23QA1400200), and Natural Science Foundation of Shanghai (23ZR1403500).
 Appendix A Languages in Evaluation Corpus
@@
 The 30 languages and their respective token counts (use LLaMA-2 Tokenizer) are as follows: Arabic (4702998), Chinese (2869208), Czech (1362041), Danish (36467), Dutch (3991305), English (1216599), Finnish (372303), French (6755281), German (2884921), Greek (474622), Hungarian (1229433), Indonesian (19226), Italian (6332560), Japanese (501899), Korean (2730794), Malay (5842), Malayalam (1489244), Norwegian (42289), Persian (1736589), Polish (4948702), Portuguese (7598161), Romanian (1381598), Russian (5205716), Spanish (7163860), Swahili (630), Swedish (1450236), Tamil (2920808), Turkish (2484186), Ukrainian (455720), Vietnamese (3606202).
-Appendix B Evaluation on 30 Languages
-The regions are localized from six languages: Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese.
+Appendix B Core Linguistic Region
+The regions are localized from six languages: Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese, respectively.
+Our work does not alter the embedding layer, as we think it equates to a mapping of tokens, which does not involve modeling linguistic competence.
+Region Visualization
+In Figure 9, we present the distribution of the ‘Top’ 5% regions of the LLaMA-2-13B model on the Attn.o matrix.
+The results indicate that across various layers, the core linguistic region on Attn.o matrix is concentrated on different rows.
+This difference is observed among the 40 attention heads in the 13B model.
 Removal 3% ratio (100K)
@@
 LLaMA-2-7B perplexity on 30 languages when the removal ratio is changed to 1% and 5% ratio, with $100,000$ equivalent samples for each language. Refer to Table 17 for more details.
-Appendix C Remove and Freeze Fine-tuning
-In Table 6, we expand removing-freezing ratio to $5\%$ for LLaMA-2-7B fine-tuning, and also observe that removing the ‘Top 5%’ regions destroys the original cross-lingual generalization competence.
-Testing Dataset (Language) # Training Samples (Chinese) Removal Ratio $=5\%$
-Top & Freeze Bottom & Freeze Top & Unfreeze
-Wechat (Chinese) 0K 204630.063 6.454 204630.063
-2K 816.94 6.057 6.05
-5K 302.806 6.039 6.054
-10K 167.891 6.285 6.315
-20K 57.355 6.525 6.56
-50K 18.92 6.154 6.173
-200K 8.8 5.582 5.584
-Falcon (English) 0K 66877.711 14.027 66877.711
-2K 248218.656 14.486 14.44
-5K 476895.25 14.831 15.074
-10K 467330.656 16.248 16.334
-20K 131648.75 18.579 18.743
-50K 25494.166 29.576 29.377
-200K 6948.663 321.353 504.645
-Table 6: Removing-freezing analysis at $5\%$ removal ratio in different regions of LLaMA-2-7B.
-‘Top/Bottom’ denotes the removal region, while ‘Freeze/Unfreeze’ indicates whether the corresponding region is frozen after removal.
-Appendix D Attn. Dimensional Removal Evaluation
+Appendix C Attention Dimensional Removal
 Figure 7 (left) illustrates that the columns of the Attn.k/q/v matrices in the attention layer, as well as the rows of the Attn.o matrix, correspond to different attention head parameters.
-Conversely, the rows of the Attn.k/q/v matrices and the columns of the Attn.o matrix are closely associated with features in the representation space.
-We remove the ‘Top’ dimensions in the attention layer, and the results is displayed in Tables 7 and 8.
-Table 7 reveals that removing the ‘Top’ dimensions continues to produce more detrimental effects than other dimensions.
+Conversely, the rows of the Attn.k/q/v matrices and the columns of the Attn.o matrix are closely associated with dimensional features in the representation space.
+We remove the ‘Top’ dimensions in the attention layer, and the results is displayed in Tables 6 and 7.
+Table 6 reveals that removing the Attention layers’ ‘Top’ dimensions continues to produce more detrimental effects than other dimensions.
 The visualizations in Figure 2 show that these dimensions are largely concentrated in a few attention heads, suggesting that some attention heads contribute more significantly to the model’s linguistic competence.
 Table 7 indicates that the removals under the second setting cause more damage than the first.
-Considering that, in the second setting, the ‘Top’ dimensions in the matrix directly interact with the corresponding features in the representational space, we can conjecture that these features are tightly linked with the model’s linguistic competence.
+Considering that, in the second setting, the ‘Top’ dimensions in the matrix directly interact with the corresponding dimensional features in the representational space, we can conjecture that these features are tightly linked with the model’s linguistic competence.
 Figure 7: One can see from the left that each row of the Attn.o ($W_{o}$) corresponds to a particular attention head, and each column of the Attn.q/k/v ($W_{q/k/v}$) matrix corresponds to one as well.
@@
 Different from Table 6, in this table, the columns of the Attn.o and the rows of the Attn.K/Q/V are removed.
-Appendix E Single Parameter Perturbation
+Appendix D Single Parameter Perturbation
 In a Transformer block, each column in the Attn.o and the MLP.down matrix of the feed-forward layer can be considered as the input weights of a neuron.
@@
 Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight.
-Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons (In Figure 6 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix).
-Appendix F Individual Language Family Region
-Tables 9-13 demonstrate LLaMA-2-7B perplexity after removing Arabic, Spanish, Chinese, Korean, and Vietnamese regions, respectively.
+Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons.
+In Figure 7 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix.
+Perturbation Parameter Perplexity
+- - 5.865
+Reset 1 L1-N2100 83224.078
+Reset 1 L1-N2800 5.860
+Reset 1 L1-N4200 5.858
+Mul 10 L1-N2100 4363.462
+Mul 10 L1-N2800 5.859
+Mul 10 L1-N4200 5.864
+Table 8: Perplexity of LLaMA-2-13B on Chinese when perturbing a single weight parameter.
+Here, ‘Reset $1$’ represents resetting the parameter to $1$ (the initial value before pre-training), ‘Mul 10’ represents multiplying the parameter by $10$.
+‘L1’ represents $1$-st layers.
+‘N’ represents the ‘Input_LayerNorm’ module, followed by the perturbed dimension.
+Appendix E Ablation Study
+Tables 9 illustrate the perplexity of LLaMA-2-7B after removing core regions with and without outlier dimensions, respectively.
+The ablation experiments reveal that different methods of disruption and varying model sizes exhibit different rates of PPL collapse:
+1) Removing according to Attention.Head (attn.k/q/v.col + attn.o.row) results in a slower collapse than according to Dimensional Features (attn.k/q/v.row + attn.o.col).
+2) The 13B model shows a slower rate of collapse.
+3) The abnormal dimension is mainly concentrated in the FFN layer of the “core linguistic region”.
+If preserving outlier dimension, the speed of PPL collapse by removing FFN layers decreases most obviously, while Attention.Head is almost unaffected.
+Removal Region $N_{d}$ LLaMA-2-7B Top(100K)
+w/ outlier d w/o outlier d
+Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) 1 848.326 27.265
+3 72594.445 57308.313
+5 48001.992 44730.059
+10 62759.516 73425.438
+Attn.o(row) Attn.k/q/v(column) 1 9.731 9.732
+3 25.82 25.822
+5 62.794 23.296
+10 875.016 860.645
+Attn.o(column) Attn.k/q/v(row) 1 167.804 9.586
+3 68554.1 136.318
+5 4259.861 688.476
+10 68170.25 431317.863
+FFN.up/gate(row) FFN.down(column) 1 20.039 6.727
+3 74905.046 7.672
+5 114725.578 9.946
+10 239015.812 16.913
+Table 9: Removing-freezing analysis at $1\%$ removal ratio in different regions of LLaMA-2-7B.
+‘Top/Bottom’ denotes the removal region, while ‘Freeze/Unfreeze’ indicates whether the corresponding region is frozen after removal.
+Appendix F Monolingual Region
+Region Visualization
+In Figure 10, we present the distribution of the Attn.q matrix for ‘Arabic’ and ‘Vietnamese’ in 4 different layers.
+The results reveal that across various layers, the two monolingual regions are concentrated in different columns of the matrix.
+Region Removal
+Tables 10-14 demonstrate LLaMA-2-7B perplexity after removing Arabic, Spanish, Chinese, Korean, and Vietnamese regions, respectively.
 The region is obtained by removing the intersections with other languages’ respective regions from the 1% ‘Top/Bottom’ regions, selected from 10,000 or 100,000 sentences during further pre-training according to Equation 4.
+Case Study
+In Figure 8, we use the prompt "There are 365 days in a year and 12" to test the model’s output in English, Arabic, and Chinese, respectively.
+The results indicate that removing the monolingual regions causes the model to lose the relative language competence, leading it to generate repetitive, nonsensical responses rather than correct answers like "12 months in a year".
+Figure 8: Model’s generation with monolingual regions removed.
+Here, we use "There are 365 days in a year and 12" as prompt, and translate it into Arabic and Chinese to evaluate model’s performance in three languages.
+Figure 9: Visualization of the linguistic competence region (the ‘Top’ 5% region) on Attention.o matrix across 8 different layers.
+The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the ‘Top’ region.
+Figure 10: Visualization of the monolingual regions for ‘Arabic’ and ‘Vietnamese’ across 4 different layers of the Attn.q matrix.
+The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the monolingual regions.
 Languages Arabic (10K) Arabic (100K)